OpenStudy (anonymous):

Let C be the curve generated by r(t) = (t,t^2,e^t^2) for 0 to 2. Evaluate the integral xzdx+2ydy+(x^2-y)dz

OpenStudy (anonymous):

first can you distinguish x y and z

OpenStudy (anonymous):

x=t, y=t^2 and z=e^t^2

OpenStudy (anonymous):

alright so using these can you figure out way of writing x, y and z with variables of x,y,z